OpenStudy (anonymous):

Simplify the expression (1-2sin^2 θ)^2 + 4sin^2 θcos ^2 θ

OpenStudy (perl):

we can do a substitution cos(2t) = 1 - 2sin^2(t)

OpenStudy (perl):

and sin(2t) = 2 sin t cos t

OpenStudy (perl):

(1-2sin^2 t)^2 + 4sin^2 t cos ^2 t =(cos 2t) ^2 + (2 sin t cos t ) ^2 =(cos 2t ) ^2 + ( sin 2t ) ^2 can you simplify it now?

OpenStudy (anonymous):

yes thanks.